Determine whether f(x) = 2x and g(x) = 0.5x are inverses of each other.
a. Yes
b. No

Answer by stanbon(48535)

Determine whether f(x) = 2x and g(x) = 0.5x are inverses of each other.
g(f(x))=g(2x)=0.5(2x)=x
f(g(x))=f(0.5x)=2(0.5x)=x
Therefore g and f are inverses of each other.
Cheers,
Stan H.